# Evaluating Human-AI Collaboration: A Review and Methodological Framework

### v1

@@ 4.2 Subfactors & Metrics of the Framework
Task allocation in HAIC focuses on strategically distributing responsibilities between humans and AI to leverage the strengths of each party and optimize overall performance. Effective task allocation enhances efficiency, minimizes errors, and increases user satisfaction, making it a critical component in the success of HAIC systems [110]. Key subfactors in Task Allocation, and its relevant metrics include:
  • Complementarity: Aligning tasks with the unique capabilities of humans and AI. For instance, AI’s analytical strengths can complement human decision making and creativity [111].
@@
    • Query Efficiency: Indicates the number of queries the AI needs to reach a certain level of accuracy, with fewer queries signifying more efficient learning [84].
@@
    • Expertise Utilization: The degree to which human expertise is leveraged, often measured qualitatively through observation or interviews, or quantitatively through metrics like the percentage of decisions where human input is decisive [26].
  • Flexibility: The ability to adjust task allocation dynamically in response to changing circumstances or new challenges is crucial for maintaining operational effectiveness [112]. Research by Dubey et al. (2020) [3] explores how AI can be used to optimize task allocation for improved efficiency.
@@
    • Dynamic Task Allocation: The ability to adjust task distribution based on real-time needs, assessed by how quickly and effectively the system adapts to changes in workload, priorities, or available resources, as explored in studies on human-robot collaboration by Tsarouchi et al. (2017) [112].
  • Efficiency: Focuses on maximizing productivity and resource utilization by optimizing task distribution, often allowing AI to automate routine tasks [111, 39]. Efficiency can be measured by:
@@
    • Model Improvement Rate: The rate of improvement in AI performance, calculated by taking the difference between the AI’s performance at a later time point and its performance at an earlier time point. This difference is then divided by the time interval between these two points. This can be applied to various AI performance metrics like accuracy or error rates, and its significance in streamlining HAIC has been explored by studies such as Vossing et al. (2022) [65].
@@
    • Resource Utilization: The efficiency in using computational resources like CPU time, memory, or network bandwidth. This can be measured as a percentage of utilized resources compared to the total available, and can be a key indicator of overall system efficiency.
@@
    • Learning Curve: The rate of improvement in AI performance over time, as defined earlier in the Goals section.
  • Responsiveness: The capability of systems to rapidly adjust to new requirements or unexpected situations is essential for maintaining workflow continuity, as discussed by Johannsmeier and Haddadin (2016) [113]. Responsiveness can be evaluated through:
@@
    • Response Time: The average time taken for the AI or human to respond to queries or feedback. A lower response time indicates quicker and more effective communication, as highlighted in studies on human-robot interaction by Hemmer et al. (2022) [111, 82].
@@
    • Real-time Performance: The ability to handle time-sensitive tasks, measured by factors like throughput (tasks completed per unit time) and latency (delay in response) [94]. This is particularly important in applications like real-time decision support systems and autonomous vehicles.
@@
    • Learning Curve: The rate of improvement in AI performance over time, as defined earlier in the Goals section.
  • Collaborative Decision Making [111]: Involves integrating the insights and expertise of both humans and AI to make informed decisions, ensuring that all parties contribute to problem- solving.
@@
    • Impact of Corrections: Measured improvement in system performance resulting from human intervention, often quantified by changes in accuracy or error rates after corrections are made. This can be used to assess the effectiveness of human input in refining AI outputs [96, 95].
@@
    • Decision Effectiveness: The impact of decisions on the overall task success, measured by comparing outcomes of decisions made with and without AI collaboration. This metric can be customized based on the specific goals of the HAIC system [98, 97].
  • Continuous Learning: AI systems should evolve by learning from past interactions to enhance future performance and adaptation. Yin et al. (2021) highlight the importance of continuous learning for optimizing task allocation in dynamic environments [114].
    • Knowledge Retention: The AI’s ability to retain learned information over time, often assessed by testing the AI’s performance on tasks after a period of non-use. This metric is important for understanding the long-term impact of human feedback and training on the AI system.
    • Model Improvement Rate: (same as in Efficiency)
    • Learning Curve: (same as in Goals)
  • Mutual Support: Promoting a supportive environment where both humans and AI can offer reciprocalassistance enhances the overall collaboration effectiveness [111]. Mutual support can be evaluated through:
@@
    • Error Reduction Rate: Measures the decrease in errors or mistakes made by the AI following human intervention, showing the effectiveness of human corrections  [4].
@@
    • Confidence: The level of user trust in AI recommendations, often measured using Likert-scale questionnaires or surveys [100]. Higher confidence scores indicate a stronger belief in the AI’s capabilities and a greater likelihood of accepting its recommendations.
@@
    • Task Completion Time: Comparison of task completion times with and without AI support [101]. This metric helps assess the efficiency gains resulting from HAIC.
  • Robustness: Ensuring the AI system performs reliably under various conditions, including adversarial attacks and different domains, is crucial for HAIC effectiveness. Robustness can be evaluated through:
@@
    • Adversarial Robustness: The resilience of the AI system to adversarial attacks, measured through testing against adversarial examples.
@@
    • System Reliability: The overall system uptime and fault tolerance, measured through system logs and reliability tests.
@@
{adjustwidth}
@@
Table 3: Evaluation Metrics for Human-AI Collaboration
@@ 4.3 Building the Framework & Quantifying the Evaluation of HAIC
Questions in this mode emphasize the AI’s autonomous capabilities and performance. We focus on efficiency metrics (like Model Improvement Rate) and uncertainty handling, as these are core challenges for AI systems operating primarily independently. The AI-Centric mode is particularly relevant for systems where AI plays a dominant role in decision-making, such as predictive analytics and automated systems.
How efficiently does the AI make decisions independently?
  • Learning Curve: Measures the rate of improvement in AI performance over time.
  • Model Improvement Rate: Measures improvement over time, indicating decision-making efficiency.
  • Query Efficiency: Evaluates how much information is gained with minimal queries.
  • Resource Utilization: Assesses how well the AI manages computing power.
  • Response Time: Tracks how quickly the AI makes decisions.
  • Real-time Performance: Evaluates the AI’s ability to handle time-sensitive tasks.
  • Objective Fulfillment Rate: Measures the percentage of tasks successfully completed.
How well does the AI handle uncertainty?
  • System Accuracy: Measures overall performance, partially reflecting uncertainty handling.
  • Error Reduction Rate: Indicates if the AI makes fewer mistakes over time.
  • Prediction Accuracy: Quantifies correct outputs specifically in uncertain scenarios.
  • Feedback Impact: Shows how much human guidance improves AI decisions under uncertainty.
@@
Here, the questions center on how well the AI communicates with and supports the human user. Metrics like Clarity of Communication and Ease of Use are essential for a positive user experience. The Human-Centric mode is particularly relevant for systems where human input is crucial for decision-making, such as in creative fields or complex problem-solving.
How clear and usable is the AI’s communication?
  • Clarity of Communication: Directly assesses if queries and responses are understandable.
  • Ease of Use: Measures overall user satisfaction with the interaction.
  • User Satisfaction: Indicates if the AI’s communication methods are acceptable to humans.
  • Frequency of Feedback: Quantifies the number of feedback instances.
How effectively does the AI support human tasks?
  • Teaching Efficiency: Gauges how much effort is needed to train the AI.
  • Error Reduction Rate: Quantifies how much the AI helps in reducing human errors.
  • Confidence: Measures user trust in the AI’s recommendations.
  • Task Completion Time: Indicates if AI assistance makes tasks faster.
@@
This mode’s questions strike a balance between adaptability and task allocation. We ask about the system’s flexibility in response to feedback (Adaptability Score) and the effectiveness of resource distribution, reflecting the collaborative core of Symbiosis.
How easily does the AI adapt to human feedback?
  • Adaptability Score: A composite metric reflecting flexibility and responsiveness.
  • Strategy Adaptation Effectiveness: Measures how well the AI changes approaches based on input.
  • Feedback Quality: Assesses how valuable the provided human feedback is.
  • Feedback Impact: Shows the improvement resulting from human guidance.
How well are tasks allocated to optimize both human and AI strengths?
  • Dynamic Task Allocation: Reflects the ability to shift task distribution in real-time.
  • Query Efficiency: Assesses how well information is requested to leverage human expertise.
  • Expertise Utilization: Measures the degree to which human skills are leveraged.
  • Knowledge Retention: Evaluates if the AI learns effectively from human input.
  • Model Improvement Rate: Measures how much human input contributes to AI progress.
  • Decision Effectiveness: Indicates the impact of decisions on overall task success.
Advantages of the Decision Tree Approach
This methodology offers several benefits:
  • User-Friendly: Simplifies the evaluation process, making it accessible to users with varying levels of expertise.
  • Tailored Assessment: Ensures that the evaluation is relevant to the specific characteristics and goals of the HAIC system.
  • Adaptable Framework: Accommodates the evolving nature of HAIC technologies and practices.
Although the decision-tree based framework offers a solid foundation for evaluating HAIC, the field continues to make significant advancements. As AI systems advance and expand their range of uses, various challenges and prospects arise. In order to maintain effectiveness, it is necessary for our evaluation methods to also adjust and change.
@@
While the decision tree and associated metrics lead the evaluation process, we recognize the importance of a weighting mechanism to provide a more comprehensive evaluation of HAIC effectiveness. This system would weight each metric according to its relevance to the specific HAIC mode and the evaluation’s criteria. For example, in an AI-Centric system, accuracy might be assigned a higher weight than user satisfaction. Conversely, in a Human-Centric system, ease of use and trust could be prioritized. In the Symbiotic mode, the weights might be distributed more evenly across metrics reflecting collaboration and mutual benefit.
The overall score for a HAIC system could then be calculated as the weighted average of the individual metric scores. However, determining appropriate weights requires further research and validation across different real - world paradigms.
Several studies in the literature have utilized weighting mechanisms to evaluate complex systems. For example, Kamar et al. (2012) [116] employed a weighted approach to combine human and machine contributions in a crowdsourcing system, emphasizing the importance of weighting in achieving balanced evaluations. Similarly, Amershi et al. (2014) [83] discussed the role of weighting different interaction metrics to assess the effectiveness of human-AI collaboration tools, highlighting how different priorities can shift the weighting schema.
Moreover, Vossing et al. (2022) [65] implemented a weighting mechanism to balance transparency and performance in AI systems, demonstrating that context-specific weighting can enhance the evaluation’s relevance and accuracy. These examples underscore the practical value of weighting mechanisms and support their inclusion in our proposed framework.
By incorporating a validated weighting mechanism, our framework ensures a nuanced and context-sensitive assessment of HAIC systems, aligning with best practices in the literature. Future research should focus on refining these weights through empirical studies across various HAIC applications to enhance the robustness and applicability of the evaluation framework.
@@ 5.1 Manufacturing
The manufacturing industry’s focus on safety, accuracy, and productivity closely corresponds with the Symbiotic mode of HAIC, which requires the collaboration of human expertise and AI skills to get the best results. This mode prioritizes metrics like Adaptability Score and Confidence. Recent study highlights a shift from focusing primarily on productivity- centered measures to a more worker-centric approach, placing greater emphasis on safety and trust. Research such as Rabby et al.’s (2020) [117] time-driven performance-aware trust model and Okamura and Yamada’s (2020)  [118] adaptive trust calibration approach exemplify this change. These works emphasize the importance of the Interaction factor, specifically focusing on the subfactors of Communication Methods and Adaptability. They achieve this by creating systems that can adapt to human actions and uncertainties. The primary emphasis of Amin et al.’s (2020)  [119] mixed-perception method is on safety, which aligns with the Task Allocation factor’s concentration on Complementarity and Mutual Support. This strategy involves the collaboration between human and AI systems to promote worker protection. The focus placed by Ming Hou (2021)  [120] on context-based interface design provides additional support for the Symbiotic method. Our methodology proposes the following essential criteria for evaluating HAIC systems in manufacturing: Adaptability Score, Error Reduction Rate, Confidence, and Task Completion Time. These measurements will assess the systems’ capacity to adapt to human operators, enhance safety, foster worker confidence, and ultimately optimize industrial processes.
@@ 5.2 Healthcare
The integration of AI in healthcare, particularly in medical imaging, often operates in a collaborative mode, where AI augments human expertise to achieve more accurate diagnoses. Although AI is highly effective in medical image classification and segmentation  [121], human skill is still essential for the interpretation and ultimate diagnosis. Contributions to this collaborative approach are made by studies such as Zheng et al.’s (2023)) [122], Eldin & Kaboudan’s (2023) [123] study of AI-driven platforms, and Aleid et al.’s (2023) [53] work on tumor identification. These developments specifically focus on the Goals aspect of the HAIC framework, with a special emphasis on enhancing diagnostic Accuracy and Efficiency. The Communication Methods subfactor is crucial as AI systems must provide unambiguous and comprehensible outputs, such as the Clarity of Communication. In the ICU setting, as explored by Dziorny et al. (2022) [124], the quality of Decision Support provided by AI becomes crucial, emphasizing the need for evaluating metrics like Decision Effectiveness. Our approach emphasizes key parameters such as System Accuracy, Prediction Accuracy, Response Time, and Clarity of Communication to successfully evaluate these collaborative HAIC systems. Future study should investigate the impact of such systems on the workflow of radiologists and oncologists, specifically in relation to Decision Support and Time-to-Diagnosis. Furthermore, conducting research to examine any possible effects on Patient Communication would be important.
@@ 5.3 Finance
The finance sector largely employs HAIC in a Symbiotic manner, leveraging AI’s analytical capabilities alongside human financial experience. This collaboration is evident in the work of Atadoga et al. (2024) [125], Al-Fatlawi et al. (2024)  [54], Jain (2023) [126], and Ali et al. (2022) [127]. These studies highlight the importance of Goals such as Accuracy and Efficiency in minimizing fraud and operational costs. However, the Symbiotic nature of this collaboration also emphasizes the Interaction factor, specifically the subfactor of Clarity of Communication, as AI-generated insights must be clearly communicated to human decision-makers.The Task Allocation factor is evident in the division of labor, with AI systems often identifying potential fraud while human experts analyze complex cases. The Error Reduction Rate is a key metric to assess the effectiveness of AI in this context. Future research should investigate the ethical and security implications of these systems, along with their long-term impacts on client relationships and the quality of financial decisions.
@@ 5.4 Education
HAIC in Education demonstrates a hybrid of Symbiotic and Human-Centric paradigms, with AI playing a supportive role in enhancing both teaching practices and learner experiences. Studies by Ifenthaler and Schumacher (2023) [128], Ji et al. (2023) [129], Wang et al. (2020)  [40], and Sharma et al. (2023) [49] highlight these modes. These works target Goals such as enhancing Teaching Efficiency and achieving personalized learning through Individualized Goals. The Interaction factor is crucial, as effective communication between AI tutors and learners impacts the learning process, necessitating assessment of Clarity of Communication and Ease of Use. The Lumilo project, investigated by Holstein and Aleven (2022) [12], shows Symbiotic collaboration for improved learning outcomes, where real-time analytics provided by AI enhance the teacher’s ability to provide Mutual Support to students. Evaluating the impact of these systems requires metrics like Task Completion Time and Learning Curve to ensure a comprehensive understanding of both student and teacher experiences. Further research should emphasize fostering student agency and critical thinking within HAIC frameworks, moving beyond passive consumption of AI-generated tools toward active engagement and co-creation of knowledge, as advocated by Fadiya (2023)  [130].
6 The Unique Challenge of Evaluating Creative and Linguistic AI
As seen in the previous sections, establishing comprehensive frameworks for assessing HAIC is vital, but certain applications pose distinct challenges due to their unique nature. Two prominent examples are Large Language Models (LLMs) and Generative AI in the Arts, where AI’s impact on human creativity, expression, and communication requires specialized evaluation methodologies.
While our existing framework provides a foundation for HAIC assessment, these creative and linguistic domains raise questions about the adequacy of traditional metrics. Can quantitative measures, such as the ones listed in Table 3 fully capture the nuances of artistic collaboration or the impact of AI on human language? How can we evaluate the ethical implications of AI- generated art or the potential for bias in LLMs? These challenges necessitate a more nuanced and tailored approach to evaluation, as we will explore in the following sections.
6.1 Large Language Models (LLMs)
Large Language Models (LLMs) have revolutionized text generation, translation, and dialogue. The GPT-3 [131] family, including ChatGPT and GPT-4, showcases the power of these models to engage in complex natural language processing [132]. While their potential for transforming Human-AI Collaboration (HAIC) is vast, evaluating these interactions requires a critical lens and the development of new methodologies [133].
Ensuring the interpretability and fairness of LLMs is crucial. Binz & Schulz’s (2022) [134] study reveals that even when LLMs perform well on cognitive tasks, they might employ reasoning patterns that are opaque or reflect biases embedded in their training data. Evaluation frameworks must incorporate methods for assessing how well LLMs explain their outputs and whether biases are mitigated over time. Additionally, the quality of human interaction with LLMs is paramount. Using GPT-3 to simulate human interactions, as in the work by Aher, Arriaga, & Kalai (2022) [56], demonstrates the value of testing how well LLMs adapt to different users and communication styles.
LLMs promise to augment cognitive tasks, enhance creativity, and personalize learning. Webb, Holyoak, & Lu (2022) [135] demonstrate an LLM’s capability for analogical reasoning, a core element of problem-solving. Further research should evaluate how such abilities impact HAIC success in real-world tasks. Regarding creativity, Douglas (2023) [136] highlights LLMs’ transformative role, inviting us to develop metrics for assessing how these tools amplify human ideation and artistic expression. In education, Liu et al. (2023) [137] explore the potential of ChatGPT and GPT-4 models to tailor learning experiences. Evaluating their effectiveness in this context will necessitate student-centered metrics focused on knowledge gains and learner autonomy.
Longitudinal studies are vital for a nuanced understanding of LLM-based HAIC. Work by Yin et al. (2023) [138] on LLM self-knowledge suggests ways to track how both the AI’s capabilities and the user’s trust in the system evolve over extended periods of interaction. The work of Dwivedi et al. (2023) [139] highlights crucial challenges posed by ChatGPT, reminding us that even with the potential benefits, addressing societal implications and ethical use is paramount.
The development of robust evaluation frameworks for LLM-powered HAIC is a rapidly evolving field. By addressing interpretability, bias, the quality of interaction, and the transformative potential of LLMs on human cognition, such frameworks will drive the creation of responsible, effective, and truly collaborative AI systems.
6.2 Generative AI in the Arts
Generative AI is revolutionizing artistic expression by enabling algorithms to generate artworks in various modalities, including visual art, music, and poetry. This technology transforms AI systems from tools into active co-creators, expanding the boundaries of what artists can achieve. Notable examples include AI-generated paintings, musical compositions, and even interactive installations tailored to audience responses. Studies like that of Gmeiner, Holstein, & Martelaro (2022) [140] suggest that understanding how effective human teams learn can provide a framework for designing generative AI systems that maximize collaborative potential.
Evaluating HAIC in the context of generative AI for artists poses several challenges. Traditional art metrics such as aesthetic quality, originality, technical skill, audience reception, market value, and art historical significance provide a starting point for this evaluation [141]. These metrics, which typically focus on the final artwork, can be adapted to also consider the collaborative process between the artist and the AI. Verheijden & Funk’s (2023) [57] work highlights the importance of assessing the generative AI’s impact on the design process itself - how it boosts inspiration, facilitates communication, and shapes the creative journey.
The field necessitates methodologies that evaluate AI-assisted artistic creation holistically. Frameworks inspired by HAIC models developed for other domains will require adaptation to consider factors unique to artistic collaboration. He, Shrestha, Puranam, & Miron-Spektor (2023) [142] propose a novel theory of co-creativity as a joint search process, offering an insightful lens for analyzing the dynamic between human artist and generative AI system.
The rise of AI-assisted art brings both ethical considerations and profound societal implications. Issues of copyright, ownership, and the meaning of authenticity in this new landscape demand attention. Epstein et al. (2023) [143] call for interdisciplinary investigation into generative AI’s cultural, economic, and legal impacts. Additionally, as AI tools become ubiquitous, reevaluating the definition of "artist" and determining the skills that foster success in this evolving domain become essential considerations.
As generative AI in art continues to mature, its transformative potential is immense. We might witness a shift towards real-time, personalized art experiences, requiring HAIC evaluation models to be responsive and adaptive. Suh, Youngblom, Terry, & Cai (2021) [144] offer intriguing insights into how AI might reshape social interactions within creative settings. Moreover, focusing on documenting the artist-AI creative process itself could become as paramount as evaluating any final product.

### v2

@@ 3.1 Subfactors & Metrics of the Framework
Task allocation in HAIC focuses on strategically distributing responsibilities between humans and AI to leverage the strengths of each party and optimize overall performance. Effective task allocation enhances efficiency, minimizes errors, and increases user satisfaction, making it a critical component in the success of HAIC systems [108].
Table 5 summarizes the key subfactors and metrics associated with task allocation, providing concise definitions and measurement methods.
Table 5: Summary of metrics for the Task Allocation factor in Human-AI Collaboration.
Subfactor Metric Definition and Measurement
@@
Complementarity Query Efficiency Indicates the number of queries the AI needs to reach a certain level of accuracy. Fewer queries signify more efficient learning [78].
@@
Expertise Utilization The degree to which human expertise is leveraged. Measured qualitatively through observation or interviews, or quantitatively through metrics like the percentage of decisions where human input is decisive [19].
@@
Flexibility Dynamic Task Allocation The ability to adjust task distribution based on real-time needs. Assessed by how quickly and effectively the system adapts to changes in workload, priorities, or resources [109].
@@
Efficiency Model Improvement Rate The rate of improvement in AI performance. Calculated as the difference in performance between two time points, divided by the time interval. Applied to metrics like accuracy or error rates [59].
@@
Resource Utilization The efficiency of computational resource use (e.g., CPU time, memory). Measured as a percentage of utilized resources compared to the total available.
@@
Learning Curve The rate of improvement in AI performance over time, as previously defined in the Goals section.
@@
Responsiveness Response Time The average time taken for AI or human participants to respond to queries or feedback. A lower response time indicates quicker and more effective communication [110, 82].
@@
Real-time Performance The ability to handle time-sensitive tasks, measured by factors like throughput (tasks completed per unit time) and latency (delay in response) [90].
@@
Learning Curve The rate of improvement in AI performance over time, as defined earlier.
@@
Collaborative Decision Making Impact of Corrections Measures the improvement in system performance from human intervention. Quantified by changes in accuracy or error rates after corrections [92, 91].
@@
Decision Effectiveness The impact of decisions on task success. Measured by comparing outcomes of decisions made with and without AI collaboration [94, 93].
Continuous Learning Knowledge Retention Assesses the AI’s ability to retain learned information over time. Evaluated by testing performance after periods of non-use [111].
Model Improvement Rate Same as defined in the Efficiency section.
Learning Curve Same as defined in the Goals section.
@@
Mutual Support Error Reduction Rate Measures the decrease in errors made by the AI following human intervention. Demonstrates the effectiveness of human corrections [2].
@@
Confidence The level of user trust in AI recommendations. Measured using Likert-scale questionnaires or surveys [105].
@@
Task Completion Time Compares task completion times with and without AI support. Helps assess efficiency gains from HAIC [97].
@@
Robustness Adversarial Robustness The resilience of the AI system to adversarial attacks. Measured through testing against adversarial examples.
@@
System Reliability The overall system uptime and fault tolerance. Evaluated through system logs and reliability tests.
@@
Table 6: Evaluation Metrics for Human-AI Collaboration
Metric Formula/Description Example
Learning Curve Graphical representation of performance vs. number of training samples. In a medical image analysis task, a graph shows the AI’s accuracy in diagnosing tumors steadily increasing over 1000 training images.
@@
[…]
Feedback Impact Impact = $\text{Performance Post-Feedback}-\text{Performance Pre-Feedback}Performance Post-Feedback - Performance Pre-Feedback$ A language translation AI’s BLEU score improves from 0.65 to 0.72 after incorporating user feedback on mistranslations.
Adaptability Score Adaptability Score = $\text{Performance Post-Adaptation}-\text{Performance Pre-Adaptation}Performance Post-Adaptation - Performance Pre-Adaptation$ A self-driving car adapts its driving style based on weather conditions, improving safety scores by 12%.
[…]
Impact of Corrections Improvement = $\text{Performance Post-Correction}-\text{Performance Pre-Correction}Performance Post-Correction - Performance Pre-Correction$ Human corrections to an AI writing assistant’s suggestions result in a 15% improvement in the text’s readability score.
[…]
Task Completion Time Time Saved = Time Without $AI\text{Time Without AI}-\text{Time With AI}Time With$out AI - Time With AI An AI-powered manufacturing robot reduces the time to complete a specific task from 10 minutes to 5 minutes.
[…]
Safety Incidents Safety Incidents = Number of safety-related issues encountered A self-driving car system logs 2 safety incidents per 1000 miles driven.
[…]
@@ 3.2 Building the Framework & Quantifying the Evaluation of HAIC
Each metric is color-coded based on its primary factor (Table 2), enhancing clarity and usability. This adaptable, scalable framework fosters a comprehensive understanding of HAIC systems, improving their assessment and development.
@@
Focuses on AI autonomy, emphasizing efficiency and uncertainty handling metrics.
@@
Evaluates AI’s communication and support for human tasks, emphasizing usability and user satisfaction.
@@
Assesses adaptability and task distribution to optimize human-AI strengths.
By structuring evaluation through this decision-tree approach, we enhance consistency, enabling better-informed HAIC system development and optimization. This approach simplifies evaluation, tailors assessments to HAIC systems, and remains adaptable as technologies evolve.
@@
While the decision tree and associated metrics lead the evaluation process, we recognize the importance of a weighting mechanism to provide a more comprehensive evaluation of HAIC effectiveness. A weighting mechanism ensures a comprehensive evaluation, adjusting metric importance based on HAIC mode. Weighted averages determine overall scores, enhancing accuracy and adaptability. Prior research supports this approach, emphasizing context-specific evaluation [113][83][59]. Future work should refine these weights through empirical studies to strengthen HAIC assessment robustness.
@@ 4 Implications of the HAIC Evaluation Framework to Real-World Domains
Although it has not yet been empirically validated, the proposed HAIC framework offers theoritical basis for assesing human-AI collaboration across sectors such as manufacturing, healthcare, finance, and education. By emphasizing factors like Adaptability, Confindence and Clarity of Communication this framework lays the groundwork for future empirical studies.
@@ 4.1 Manufacturing
Manufacturing’s focus on safety, accuracy, and productivity aligns with the HAIC Symbiotic mode, which integrates human expertise with AI skills. Studies [114, 16, 115, 116] demonstrate a shift towards worker-centric, adaptive approaches that highlight critical metrics such as Adaptability Score, Error Reduction Rate, Confidence, and Task Completion Time.
@@ 4.2 Healthcare
In medical imaging diagnosis, AI supports human decision-making by enhancing diagnostic accuracy and efficiency [117, 118, 119, 47, 120]. The framework targets improvements in diagnostic Accuracy and Efficiency, emphasizing clear AI outputs. Evaluation parameters include System Accuracy, Prediction Accuracy, Response Time, and Decision Effectiveness, with further research needed on workflow impacts and patient communication.
@@ 4.3 Finance
The finance sector benefits from a Symbiotic approach where AI’s analytical capabilities complement human expertise to enhance fraud detection and operational efficiency [121, 48, 122, 123]. Key aspects include achieving goals of Accuracy and Efficiency and ensuring clear communication of AI insights. Metrics such as the Error Reduction Rate help assess system effectiveness, with future work needed to address ethical and long-term impacts.
@@ 4.4 Education
In education, a mix of Symbiotic and Human-Centric models enhances teaching and learning [124, 125, 34, 43, 10, 126]. The focus is on improving Teaching Efficiency and achieving personalized learning through effective communication. Important metrics include Task Completion Time and Learning Curve, while further research should explore fostering active student engagement and critical thinking.